# Cislunar Constellation Design for Space Situational Awareness with Time-Expanded Facility Location Problem

### v1

@@ 6.1 Comparison of Lagrangean Method against Gurobi
$p$ Gurobi (TimeLimit $=500$) Gurobi (TimeLimit $=3600$) Lagrangean Method
@@
2 0.1061 89 OPTIMAL - - - 0.1033 540 $-2.68$ $-2.68$
3 0.1540 500 TIME_LIMIT 0.1540 3522 OPTIMAL 0.1318 520 $-14.39$ $-14.39$
4 0.1999 500 TIME_LIMIT 0.1999 3603 TIME_LIMIT 0.1837 504 $-8.09$ $-8.09$
5 0.2435 502 TIME_LIMIT 0.2435 3600 TIME_LIMIT 0.2176 535 $-10.61$ $-10.61$
@@
$p$ Gurobi (TimeLimit $=500$) Gurobi (TimeLimit $=3600$) Lagrangean Method
@@
2 0.1098 501 TIME_LIMIT 0.7259 3607 TIME_LIMIT 0.6194 530 $+463.94$ $-14.67$
3 0.1358 501 TIME_LIMIT 0.8315 3601 TIME_LIMIT 0.7334 526 $+439.96$ $-11.80$
4 0.1439 500 TIME_LIMIT 0.6670 3601 TIME_LIMIT 0.8192 513 $+469.09$ $+22.81$
5 0.1669 500 TIME_LIMIT 0.7498 3603 TIME_LIMIT 0.8454 559 $+406.39$ $+12.75$
@@
$p$ Gurobi (TimeLimit $=500$) Gurobi (TimeLimit $=3600$) Lagrangean Method
@@
2 0.1321 506 TIME_LIMIT 0.1321 3611 TIME_LIMIT 0.6934 532 $+424.92$ $+424.92$
3 0.1612 506 TIME_LIMIT 0.1612 3606 TIME_LIMIT 0.7297 529 $+352.76$ $+352.76$
4 0.1925 505 TIME_LIMIT 0.1925 3613 TIME_LIMIT 0.8080 517 $+319.81$ $+319.81$
5 0.2140 506 TIME_LIMIT 0.2140 3613 TIME_LIMIT 0.9322 511 $+335.53$ $+335.53$
Table 8: TE-$p$-MP instance with LET demand and FOV = $60^{\circ}$ solved using Gurobi (MIPGap 1%) versus Lagrangean method. All reported times are in seconds.
(a) $\bar{m}_{\mathrm{crit}}=15$
$p$ Gurobi (TimeLimit $=500$) Gurobi (TimeLimit $=3600$) Lagrangean Method
2 0.3737 39 OPTIMAL - - - 0.3144 534 $-15.86$ $-15.86$
3 0.2406 500 TIME_LIMIT 0.5078 1291 OPTIMAL 0.3967 529 $+64.89$ $-21.88$
4 0.1703 500 TIME_LIMIT 0.5959 3508 OPTIMAL 0.4681 533 $+174.92$ $-21.45$
5 0.1840 500 TIME_LIMIT 0.6674 3600 TIME_LIMIT 0.5502 545 $+199.05$ $-17.56$
(b) $\bar{m}_{\mathrm{crit}}=18$
$p$ Gurobi (TimeLimit $=500$) Gurobi (TimeLimit $=3600$) Lagrangean Method
2 0.9212 500 TIME_LIMIT 0.9398 3600 TIME_LIMIT 0.8353 524 $-9.33$ $-11.12$
3 0.4414 500 TIME_LIMIT 0.9546 3650 TIME_LIMIT 0.9056 517 $+105.16$ $-5.14$
4 0.4582 500 TIME_LIMIT 0.4582 3600 TIME_LIMIT 0.9423 586 $+105.64$ $+105.64$
5 0.5335 501 TIME_LIMIT 0.9921 2220 OPTIMAL 0.9545 500 $+78.89$ $-3.79$
(c) $\bar{m}_{\mathrm{crit}}=20$
$p$ Gurobi (TimeLimit $=500$) Gurobi (TimeLimit $=3600$) Lagrangean Method
2 0.4607 500 TIME_LIMIT 0.9701 3600 TIME_LIMIT 0.8807 615 $+91.17$ $-9.22$
3 0.5431 502 TIME_LIMIT 0.5431 3600 TIME_LIMIT 0.9461 512 $+74.21$ $+74.21$
4 0.0000 500 TIME_LIMIT 0.9937 723 OPTIMAL 0.9868 621 $+Inf$ $-0.69$
5 0.0000 502 TIME_LIMIT 0.9958 1327 OPTIMAL 0.9774 538 $+Inf$ $-1.85$
@@ 6.2 Analysis of CSSA Coverage
As a demonstration of the usability of the proposed problem formulation and the LM, we now look in detail at the trade space for designing CSSA constellations for the SOI, cone of shame, and LET transit window demands. For consistency across all cases, all results are based on solutions obtained using the LM, with a time limit of 1000 seconds. We provide insight on a select number of architectures across a few combinations of parameters. The full list of LPOs used by each problem from Gurobi and LM are provided in Tables in Appendix 7.2.
6.2.1 Sphere of Influence
Competitive solutions with the SOI demand are found to utilize planar LPOs, particularly the 1:1 resonant L1 and L2 Lyapunov orbits, and DROs of various resonances. We particularly focus on two types of solutions: the former is an example using the Lyapunov orbits, with $\bar{m}_{\mathrm{crit}}=20$ and $p=2$, and the latter is an example using the DROs, with $\bar{m}_{\mathrm{crit}}=20$ and $p=5$, shown in Figure 10.
The Lyapunov orbit-based constellation in Figure 10(a) contains two observers on the 1:1 resonant L1 Lyapunov, which extend beyond the region spanned by the SOI along the $y$-axis in the Earth-Moon rotating frame. Clock-wise rotating LPOs with 1:1 resonance are particularly well suited for optical observations since the solar phase angle $\phi_{S}$ in the direction of a static target remains nearly constant [2, 13]; while the 1:1 L1 Lyapunov has a varying distance from the set of targets, choosing the appropriate phase along the LPO provides consistently low $\bar{m}_{\mathrm{target}}$, as $\bar{m}_{\mathrm{target}}$ has a stronger dependence on $\phi_{S}$ over the distance, as shown in Figure 5. Meanwhile, the two L2 Lyapunov orbits are placed with some along-track separation, but not at $180^{\circ}$ apart from one another; this is a manifestation of the property of the clock-wise rotating 1:1 resonant LPO, where both observers are preferentially placed in angular locations that provide sufficiently good illumination conditions.
This is contrasted by the latter solution shown in Figure 10(b), where five observers are all placed in a 3:2 resonant DRO, with equal angular separation. Because the resonance is not 1:1, angular locations along the LPO with good or poor illumination conditions vary over time. Thus, the five observers are used in a complementary manner, in such a way that there are always at least two to three observers that can monitor the set of targets.
(a) Gurobi solution with $\bar{m}_{\mathrm{crit}}=20$, $p=2$
(b) Gurobi solution with $\bar{m}_{\mathrm{crit}}=20$, $p=5$
Figure 10: Select architecture solutions for Sphere of Influence demand at time-step $t=1$
@@ 6.2.2 Cone of Shame
In the Cone of Shame demand, the 1:1 resonant L1 Lyapunov orbit is found to be used in many competitive instances. Figure 11 shows the LM-based solutions for $p=3,4,5$ and $\bar{m}_{\mathrm{crit}}=20$. In these three solutions, as the number of observers $p$ is increased, an additional observer is added to the same orbit, resulting in an increase in coverage fraction as well. As it was noted from the solution in Figure 10(a), the clock-wise rotating 1:1 resonant LPOs are placed in a leading/trailing configuration along the LPO rather than being placed at equal anomaly separation. The persistence of this train-like configuration across multiple problem instances highlights the efficacy of this constellation design strategy, particularly for monitoring targets in the Cone of Shame.
(a) Lagrangean method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=3$
(b) Lagrangean method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=4$
(c) Lagrangean method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=5$
Figure 11: Select architecture solutions for Cone of Shame demand at time-step $t=1$
@@ 6.2.3 Low-Energy Transfer Transit Window
We finally look into solutions for the LET transit demand. We first compare solutions with increasing cut-off magnitude from 15 to 20, with observers fixed to $p=4$; the corresponding architectures, obtained by the LM, are shown in Figures 12(a)12(b), and 12(c). When the cut-off magnitude is lower, targets become harder to observe, and the observers are placed in a 2:1 resonant L2 Halo orbit that is closer to the set of targets; this comes with the disadvantage that the observer is able to fit fewer targets within its FOV at a given time. Meanwhile, the solutions at $\bar{m}_{\mathrm{crit}}=18,20$ use a 3:1 resonant L2 Halo orbit that is further away, offering larger coverage from a single sensor pointing allocation.
We also recognize that the train-like constellation, reminiscent of solutions from Figures 10(a) or 11, is used on the 2:1 resonant Halo for $\bar{m}_{\mathrm{crit}}=15$, and for three of the four observers on the 3:1 resonant Halo for $\bar{m}_{\mathrm{crit}}=18,20$. The fourth observer in the last two cases is placed with a significant angular separation along the LPO to complement the coverage; note that because the LPO is not planar, the illumination conditions are not necessarily poor for this last observer during the times along the LPO where there is a non-negligible out-of-plane component.
As the observer is increased to $p=5$, the architecture consists of four 3:1 L2 Halo observers and an additional observer on the 1:1 L1 Lyapunov, as shown in Figure 12(d). Now, all four observers on the 3:1 L2 Halo form a train-like configuration, all in adjacent angular locations, and a complementing observer is now placed on a different LPO altogether. Note that the 1:1 resonant L1 Lyapunov is a favorable LPO for placing a single additional observer due to the invariance of the illumination condition; while the larger orbit results in an increased distance from the target set, this does not constitute a penalty in this configuration, where $\bar{m}_{\mathrm{crit}}$ is set high.
(a) Lagrangean method solution with $\bar{m}_{\mathrm{crit}}=15$, $p=4$
(b) Lagrangean method solution with $\bar{m}_{\mathrm{crit}}=18$, $p=4$
(c) Lagrangean method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=4$
@@
Figure 12: Select architecture solutions for LET transit demand at time-step $t=1$
@@ 7 Conclusions
In this work, the TE-$p$-MP has been formulated to design satellite constellations for CSSA. The problem consists of determining the $p$ optimal observer locations along with the orientation of their optical sensors over time to observe a given set of static or dynamic targets. The resulting optimization problem is a binary linear program (BLP), which is discretized in terms of time, candidate locations of observer spacecraft, and sensor pointing directions. A custom LM based on the relaxation of complicating constraints, coupled with custom heuristics leveraging the problem structure as well as the geometry of the candidate facility locations, has been proposed.
The TE-$p$-MP, along with the LM, has been demonstrated to find near-optimal solutions for various CSSA scenarios, varying in terms of the distribution of demands or observer parameters. The primary benefit of the LM is the reduction in solve time compared to using generic BLP solvers, which enables exploring the trade space of parameters such as the visible magnitude threshold or the temporal and spatial distribution of observation targets. The LM has been demonstrated to perform in a largely consistent manner across most TE-$p$-MP instances and to deliver competitive near-optimal constellations, whereas Gurobi’s performance is found to be far more precarious, depending on the specific parameters of the instance. Several noteworthy constellations tailored to monitoring demands based on the lunar sphere of influence, the Cone of Shame, and a transit window of LETs have been analyzed to demonstrate the proposed approach. One notable remark is on the existence of train-like constellations, where observers are packed in an along-track formation; this type of solution has been found frequently when the observers are located on clockwise-rotating 1:1 resonant LPOs, where the illumination condition is nearly invariant.

### v2

@@ 6.1 Comparison of Lagrangian Method against Branch-and-Bound
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
@@
2 0.3737 66 OPTIMAL - - - 0.3144 523 $-15.86$ $-15.86$
3 0.3246 501 TIME_LIMIT 0.5078 1426 OPTIMAL 0.3967 644 $+22.20$ $-21.88$
4 0.1703 501 TIME_LIMIT 0.5973 1846 OPTIMAL 0.4681 519 $+174.92$ $-21.63$
5 0.1845 500 TIME_LIMIT 0.6670 3601 TIME_LIMIT 0.5471 537 $+196.50$ $-17.97$
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
@@
2 0.9309 500 TIME_LIMIT 0.9398 2263 OPTIMAL 0.8339 638 $-10.42$ $-11.27$
3 0.0000 501 TIME_LIMIT 0.0000 3600 TIME_LIMIT 0.9056 634 $+Inf$ $+Inf$
4 0.0000 500 TIME_LIMIT 0.0000 3600 TIME_LIMIT 0.9355 524 $+Inf$ $+Inf$
5 0.0000 500 TIME_LIMIT 0.0000 3600 TIME_LIMIT 0.9461 569 $+Inf$ $+Inf$
@@
$p$ B&B (TimeLimit $=500$) B&B (TimeLimit $=3600$) Lagrangian Method
@@
2 0.4778 500 TIME_LIMIT 0.9701 3601 TIME_LIMIT 0.8681 526 $+81.70$ $-10.51$
3 0.4780 501 TIME_LIMIT 0.4780 3600 TIME_LIMIT 0.9461 537 $+97.91$ $+97.91$
4 0.4794 500 TIME_LIMIT 0.9957 819 OPTIMAL 0.9868 547 $+105.84$ $-0.89$
5 0.4794 500 TIME_LIMIT 0.9993 551 OPTIMAL 0.9774 560 $+103.88$ $-2.20$
@@ 6.2 Analysis of CSSA Coverage
We now look in detail at the trade space for designing CSSA constellations for the cone of shame, and LET transit window demands. For consistency across all cases, all results are based on solutions obtained using the LM, with a time limit of 1000 seconds. We provide insight on a select number of architectures across a few combinations of parameters. The full list of LPOs used by each problem from Gurobi and LM are provided in Tables in Appendix 7.2.
@@ 6.2.1 Cone of Shame
Figure 10 shows the LM-based solutions for $p=2,3,4,5$ and $\bar{m}_{\mathrm{crit}}=20$. In the Cone of Shame demand, the 1:1 resonant L1 Lyapunov orbit is found to be used in many competitive instances. As an exception, when only two observers are placed, an alternative constellation configuration involving also a 3:1 resonant halo orbit is found. In these three solutions, as the number of observers $p$ is increased, an additional observer is added to the same orbit, increasing coverage fraction as well.
Clock-wise rotating LPOs with 1:1 resonance are particularly well suited for optical observations since the solar phase angle $\phi_{S}$ in the direction of a static target remains nearly constant [2, 13]. Even though the 1:1 L1 Lyapunov has a varying distance from the set of targets, choosing the appropriate phase along the LPO provides consistently low $\bar{m}_{\mathrm{target}}$, due to the strong dependence of $\bar{m}_{\mathrm{target}}$ on $\phi_{S}$ over the distance, as shown in Figure 5. To benefit from favorable constellations, we find solutions across values of $p$ and $\bar{m}_{\mathrm{crit}}$ where observers are thus placed on the 1:1 resonant Lyapunov with a leading/trailing configuration rather than being placed at equal anomaly separation. The persistence of this train-like configuration across multiple problem instances highlights the efficacy of this constellation design strategy, particularly for monitoring targets in the Cone of Shame. We note that these constellations are unlike typical Earth-bound constellations that are equally spaced along a given orbit; rather than being equally distributed along track, they form a leading/trailing formation.
The steering history for the selected architectures in Figure 10 are shown in Figure 11. For $p=2$, as shown in Figure 11(a), observer 1 located on the 3:1 resonant Halo steers its sensor both along the azimuth and elevation angles, while observer 2 has its sensor lied on the $xy$-plane. The azimuth steering angle of the two observers exhibit periodicities according to their orbital periods. We note that observer 1 is not disadvantaged by the fact that that it is not on a 1:1 resonance as its sensor is pointed out of plane with a non-zero elevation angle for the majority of time. To provide physical intuition, we plot the steering history of the $p=2$ architecture in position space in Figure 12. For the sake of clarity, we only show the first 10 steps, corresponding to one revolution along observer 1 located on a 3:1 resonant LPO. With architectures for $p=3,4,5$, we are able to see steering histories of observers on the same, 1:1 resonant L1 Lyapunov. The azimuth angle history for each observer follows a similar track shifted according to their location along the LPO. The appropriate azimuth angle to steer the sensor on the 1:1 resonant LPO is largely dictated by the Sun’s illumination direction, causing the observers to follow a similar in-plane steering strategy. Meanwhile, the elevation steering history is distinct for each observer; the out-of-plane sensor pointing direction is varied in order to achieve a more comprehensive coverage of the volume of interest.
The sensor steering history provides insights as to why in the $p=2$ case, the solution consists of one observer on the L1 Lyapunov orbit and another on a different LPO, while in cases with $p=3,4,5$ all observers placed on the L1 Lyapunov orbit. With at least $p=3$ observers, each can be tasked to observe at elevation angles of $0^{\circ}$, $-30^{\circ}$, and $30^{\circ}$ thereby complementing the observation to provide holistic coverage. Meanwhile, with only $p=2$ observers, the 3:1 resonant Halo, with higher cadence steering across azimuth and elevation angles, is found to be able to better complement the observation of a single L1 Lyapunov orbit observer. We note that the elevation angle spread for $p=3$ as seen in Figure 10(b) is not as thorough as with cases with $p=4$ and $5$, as seen in Figures 10(c) and 10(d). This is the reason why the LM solution is $-15.79\%$ worse than the B&B solution after 3600 seconds, as reported in Table 5(c), compared to cases with $p=2,4,5$.
(a) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=2$
(b) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=3$
(c) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=4$
(d) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=5$
Figure 10: Optimized observer locations for Cone of Shame demand at time-step $t=1$ (at a new Moon). Red shaded volume represents region of interest.
(a) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=2$
(b) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=3$
(c) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=4$
(d) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=5$
Figure 11: Optimized observer pointing direction histories for Cone of Shame demand
(a) Observer 1 on 3:1 resonant Northern L2 Halo
(b) Observer 2 on 1:1 resonant L1 Lyapunov
Figure 12: Sensor pointing directions of each observer over 10 time-steps for $\bar{m}_{\rm crit}=20$ $p=2$. Red shaded volume represents the region of interest.
@@ 6.2.2 Low-Energy Transfer Transit Window
We first compare solutions with increasing cut-off magnitude from 15 to 20, with observers fixed to $p=4$; the corresponding architectures, obtained by the LM, are shown in Figures 13(a)13(b), and 13(c). When the cut-off magnitude is lower, targets become harder to observe, and the observers are placed in a 2:1 resonant L2 Halo orbit that is closer to the set of targets; this comes with the disadvantage that the observer is able to fit fewer targets within its FOV at a given time. Meanwhile, the solutions at $\bar{m}_{\mathrm{crit}}=18,20$ use a 3:1 resonant L2 Halo orbit that is further away, offering larger coverage from a single sensor pointing allocation.
We also recognize that the train-like constellation, reminiscent of solutions from Figure 10, is used on the 2:1 resonant Halo for $\bar{m}_{\mathrm{crit}}=15$, and for three of the four observers on the 3:1 resonant Halo for $\bar{m}_{\mathrm{crit}}=18,20$. The fourth observer in the last two cases is placed with a significant angular separation along the LPO to complement the coverage; note that because the LPO is not planar, the illumination conditions are not necessarily poor for this last observer during the times along the LPO where there is a non-negligible out-of-plane component.
With $p=5$, the architecture consists of four 3:1 L2 Halo observers and an additional observer on the 1:1 L1 Lyapunov, as shown in Figure 13(d). Now, all four observers on the 3:1 L2 Halo form a train-like configuration, all in adjacent angular locations, and a complementing observer is now placed on a different LPO altogether. Note that the 1:1 resonant L1 Lyapunov is a favorable LPO for placing a single additional observer due to the invariance of the illumination condition; while the larger orbit results in an increased distance from the target set, this does not constitute a penalty in this configuration, where $\bar{m}_{\mathrm{crit}}$ is set high.
Figure 14 shows the sensor steering history for architectures in Figure 13. In all cases, all observers are located on one side of the region of interest; as a result, the azimuth angle remains mostly between $-30^{\circ}$ and $30^{\circ}$. With the $p=2$ case utilizing Southern halos, the elevation angle across observers and across time is skewed on the positive side, while with the $p=3,4,5$ cases utilizing Northern halos, the elevation angle is skewed on the negative side. Figure 15 shows the sensor pointing directions for $\bar{m}_{\rm crit}=20$ and $p=4$ from Figure 13(c). As the Northern halo has a larger out-of-plane amplitude than the region of interest, sensors are pointed downward around the apolunes along the Northern halo. We also note that over the first 10 time-steps, the steering directions of each observer have slight variations between one another due to the variation in their respective phase along the LPO.
(a) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=15$, $p=4$
(b) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=18$, $p=4$
(c) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=4$
@@
Figure 13: Optimized observer locations for LET transit demand at time-step $t=1$ (at a new Moon). Red shaded volume represents region of interest.
(a) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=3$
(b) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=4$
(c) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=5$
(d) Lagrangian method solution with $\bar{m}_{\mathrm{crit}}=20$, $p=5$
Figure 14: Optimized observer pointing direction histories for LET transit demand
(a) Observer 1 on 3:1 resonant Northern L2 Halo
(b) Observer 2 on 3:1 resonant Northern L2 Halo
(c) Observer 3 on 3:1 resonant Northern L2 Halo
(d) Observer 4 on 3:1 resonant Northern L2 Halo
Figure 15: Sensor pointing directions of each observer over 10 time-steps for $\bar{m}_{\rm crit}=20$ $p=4$. Red shaded volume represents the region of interest.
6.3 Key Takeaways on the Architecture Design Space
Taking the analyses on the obtained architectures into account, the following general remarks can be made:
  • For a set of demand spread across L1 and L2, the large 1:1 resonant L1 Lyapunov offers an advantageous observation location; when the sensor is insufficient to observe targets at long ranges, the solution favors an alternative LPO that resides closer to certain regions within the volume of interest where most targets reside.
  • Due to the solar phase angle largely dictating the visibility of targets, observers on 1:1 resonant LPOs tend to be in a leading/trailing configuration rather than being evenly spaced along the LPO. As expected, the pointing directions are chosen in a manner that each observer complements the volume coverage of one another.
  • For a set of demand centered around L2, LPOs covering a smaller spatial region, in particular 2:1 and 3:1 resonant L2 halos, are favored. These LPOs are suitable as the observer fly in region near but outside the volume of interest; the latter consideration is crucial when considering FOV, as residing inside the volume of interest would severely limit the fraction of volume that is visible to a given observer, despite its proximity to targets.
@@ 7 Conclusions
In this work, the TE-$p$-MP has been proposed for CSSA. The formulation consists of determining the $p$ optimal observer locations along with the orientation of their optical sensors over time to observe a given set of static or dynamic targets. The TE-$p$-MP, inspired from the classical $p$-Median problem, is a BLP discretized in time, candidate locations of observer spacecraft, and sensor pointing directions. This formulation simultaneously solves for the constellation design and the sensor-tasking problems. The formulation has been developed together with a custom LM algorithm, based on the relaxation of complicating constraints and custom heuristics leveraging the problem structure as well as the geometry of the candidate facility locations. A solution to the TE-$p$-MP is a CSSA constellation together with sensor-tasking for each observer that observes a prescribed set of targets.
The TE-$p$-MP, along with the LM, has been demonstrated to find near-optimal solutions for various CSSA scenarios varying in terms of the distribution of demands or observer parameters in the order of minutes. The primary benefit of the LM is the reduction in solve time compared to using generic BLP solvers, which enables exploring the trade space of parameters such as the visible magnitude threshold or the temporal and spatial distribution of observation targets. The LM has been demonstrated to perform in a largely consistent manner across most TE-$p$-MP instances and to deliver competitive near-optimal constellations, whereas B&B’s performance is found to be far more precarious, depending on the specific parameters of the instance. Several noteworthy constellations tailored to monitoring demands based on the Cone of Shame and a dynamic transit window of LETs have been analyzed to demonstrate the proposed approach.